# Exotic Particles at the DUNE Near Detector from Charged Pion Scattering

### v1

@@ IV.2 MCP Detector Signatures and Backgrounds
The proposed MCP detector design is based on that of MilliQan [22, 23]. The signature for detecting a MCP is three soft ionizations through the scintillator arrays that are collinear with the beam line and target, detectable through the collection of photoelectrons by PMTs. Following Ref. [2], we also require that all three soft ionizations occur within a small 15 nanosecond window to further reduce background.
There are two types of background that would look like additional soft ionizations throughout the scintillators: detector-related background and beam-related background. To make a direct comparison with Ref. [2], we make identical assumptions about the backgrounds, yielding an estimate of about 600 events per year; we refer the reader to Ref. [2] for more details about the background processes.
@@ IV.3 MCP Sensitivity at Dune ND
In the FerMINI design [2, 24], a minimum-ionizing particle with charge $|Q|=1$ would produce approximately $10^{6}$ photoelectrons (PE) in a 1 m plastic scintillator with a density of $\rho_{\rm scint.}=1$ g/cm3. Thus MCPs with charge $Q=e\epsilon$ would produce PEs that scale as $\epsilon^{2}$. We can then quantify the probability of a MCP producing at least one PE in all three scintillator stacks as
@@
where $\overline{N}_{\rm PE}$ denotes the average number of photoelectrons produced from soft ionization. We use the approximation of $\overline{N}_{\rm PE}(\epsilon)=\left(\frac{\epsilon}{\zeta}\right)^{2}$ where $\zeta=2\times 10^{-3}$. This value of $\zeta$ corresponds to a Saint-Gobain BC-408 plastic scintillator [2, 23]. Given these choices, such an experiment could be sensitive to $\epsilon$ of order $10^{-4}$.
@@
Assuming that the scintillators can detect values as low as $\epsilon=10^{-4}$, we compute the projected 95% CL limit, i.e. $N_{S}/\sqrt{N_{B}}=2$ where $N_{B}$ is the number of background events, for the DUNE ND setup to MCP from charged pion scattering. The results are shown in Fig. 6 along with constraints and projections from other beam-dump, collider, and direct detection experiments. We also show the projected limit from neutral meson decay and Drell-Yan scattering at DUNE, following Fig. 1 of Ref. [2] for the total number of MCPs that reach the target, and then applying Eq. (12) in order to model the scintillator response. We find that we reproduce the sensitivity in Ref. [2] to within about a factor of 2, with any differences likely due to slightly different treatments of the angular acceptance [25], which lets us directly assess the increase in sensitivity from charged pion scattering. Overall, the charged pion production channel (which is irreducible in this model) can offer additional sensitivity to detecting MCPs for masses $1.5\ {\rm GeV}\lesssim m_{\chi}\lesssim 5\ {\rm GeV}$.
Figure 6: 95% CL sensitivity to MCP detection at DUNE ND with $10^{21}$ POT (about one year of beam time). The charged pion production channel is shown in purple, and the red-orange curve labeled “Neutral Meson + Drell Yan” is our reproduction of the projected sensitivity from Ref. [2]. The milliQan HL-LHC projected sensitivity [22, 23] is plotted in teal, and the LANSCE-mQ at ER1 [26] projected sensitivity, using $5.9\times 10^{22}$ POT (about 3 years of beam time), is plotted in blue. Shaded regions are existing limits from SLAC ([27], peach), SENSEI ([28], yellow), BEBC ([29], light pink), Charm II ([30], light green), LSND ([31], olive green), MiniBooNE ([31], gray), ArgoNeuT ([32], light blue), and colliders ([33], burgundy).
@@ V Heavy Axion-like Particles
The DUNE ND is an excellent experiment to look for long-lived exotic particles at MeV-GeV mass scales. One such particle that we can search for at DUNE ND is a heavy axion coupled to gluons at low energies. This particle may be a generic ALP, but in some models with additional gauge groups, such a heavy axion may in fact be the QCD axion which solves the strong CP problem [34]. In this section, we will be agnostic as to the UV interpretation of such a particle, and simply use this model as a benchmark to make a direct comparison with the model of Ref. [4]. As we have done with MCPs in Sec. IV, we will demonstrate the additional sensitivity gained from charged pion production, which is an irreducible production channel in this model.
As in Ref. [4], we consider a low-energy ALP model where at scales $\mu=\mathcal{O}(1)$ GeV the ALP is only coupled to gluons,
@@
Here, $f_{a}$ is the axion decay constant, $\alpha_{s}$ is the strong fine structure constant, and $G^{a}_{\mu\nu}$ is the gluon field strength tensor. Following [35, 36, 37], the axion couplings to nucleons and pions in the chiral Lagrangian can be derived by performing a chiral rotation to shift the axion field into the quark mass matrix. This rotation is arbitrary, but in the conventional parameterization which eliminates mass mixing between the axion and the $\pi^{0}$, we obtain
@@
As in Sec. IV, $f_{\pi}=93$ MeV is the pion decay constant, $g_{A}=1.27$ is the nucleon axial coupling, and $g_{0}=0.52$ is another axial coupling extracted from lattice QCD simulations and evaluated at an $\overline{\rm MS}$ scale $\mu=2$ GeV [38]. Note that the chiral rotation which generates these couplings will also generate an axion-photon coupling; in the spirit of this “gluon dominance” model, we will ignore production from this operator and only use it to permit decays $a\to\gamma\gamma$ for $m_{a}<3m_{\pi}$ (see Sec. V.2 below). Furthermore, we note that axion-neutral meson kinetic mixing will also be present in the chiral Lagrangian. In order to compare the parametric dependence of scattering processes to previously-considered mixing processes, we are neglecting diagrams which involve kinetic mixing insertions, and we leave to future work a full analysis which simultaneously treats mixing and scattering in order to capture any potential interference effects.
@@ V.1 ALP Flux
Figure 8: The cross section for the scattering process $\pi N\to\pi Na$ as a function of incoming pion energy $E_{\pi}$ for range of axion masses and $f_{a}=10^{4}$ GeV. The cross section scales with increasing pion energy due to the derivative couplings in the chiral Lagrangian. The cross section also has a weak dependence on $m_{a}$. The $m_{a}$ dependence contributes mostly at low energies.
@@
We can understand this enhancement through a dimensional analysis estimate. The number of axions produced through neutral pion mixing is [4]
@@
This is reasonably close to the $10^{6}$ enhancement seen in Fig. 9, which accounts for all of the order-1 coupling factors, the true pion energy distribution, and the detector’s geometric acceptance.
@@ V.2 ALP Detector Signatures and Backgrounds
The axions that are created within the angular acceptance of ArgonCube and MPD at DUNE ND can be detected by their decay into SM particles. For the axion masses we are interested in, the axion can decay into photons or hadrons. Even in the gluon-dominated model, axions are able to decay to photons through mixing with neutral mesons (indeed, this was the primary production mechanism of Ref. [4]). For $m_{a}<3m_{\pi}$, the axion will primarily decay into two photons, $a\to\gamma\gamma$. The decay width is given by [39]
@@
For $3m_{\pi}<m_{a}\lesssim 1$ GeV, the axion decays to primarily to hadrons through $a\to 3\pi$ and $a\to\pi\pi\gamma$ [39, 4].
We consider the case where the axion decays within the liquid and gaseous argon detectors at DUNE ND, between 574 and 584 m from the target, with an angular acceptance of $\sim$ 3 mrad (see Fig. 1). Approximately 1% of the total number of axions produced are detected due to the small angular acceptance. These axions must also be highly boosted in order to travel 574 m and be within the angular resolution of the detectors. As a result, the decay products are also highly boosted compared to background events [4]. Much of the background in both detectors comes from neutrino-related scattering processes in the argon tanks [12, 4]. Both the diphoton and hadronic backgrounds would have low energies and would result in a roughly isotropic signature. Since our axion signal is highly boosted, we can easily distinguish signal from background events [4].
@@ V.3 ALP Sensitivity at DUNE ND
Combining the expected signal from $a\to\gamma\gamma$ and $a\to$ hadrons, we can determine the axion parameter space to which DUNE ND is sensitive. We calculate the number of signal events according to the methods described in Sec. III.3, using the axion decay length $c\tau_{a}$ from Refs. [4, 40]. The number of axion signal events is then
@@
where $N_{a}=A\Phi_{a}$ with $\Phi_{a}$ in (17), and $\text{BR}(a\to X)$ is the branching ratio of the axion decaying to either photons or hadrons [39, 40].
In Figure 10, we show the sensitivity of DUNE ND to 3 or more signal events in 10 years of data collection. The increased production rate from charged pion scattering away from meson resonances improves the projected sensitivity of DUNE in both the long-lifetime (small $1/f_{a}$) and short-lifetime (large $1/f_{a}$) regimes. Because we only calculated $N_{a}$ up to the chiral cutoff $|t|\sim(4\pi f_{\pi})^{2}$, we do not find additional sensitivity beyond the largest $m_{a}$ considered in Ref. [4]. For axions below the kaon mass, the sensitivity from charged pion scattering alone is within an order-1 factor of axions produced from kaon decay, recently argued in Ref. [5] to be the dominant production channel in this mass range. Including charged pion scattering beyond the first interaction length of the dump could yield additional sensitivity, possibly making these two production modes comparable.
Figure 10: The sensitivity of DUNE ND to 3 or more heavy ALP signal events over 10 years of data collection from charged pion scattering (purple) along with other projections and constraints. We show projections for DUNE ND (solid) along with FASER [41] (green dot-dashed). The projection for an axion signal from neutral meson and gluon-gluon fusion production are shown in blue [4]. The pink curve, adapted from Ref. [5] and rescaled to the same event rate as the other curves, shows diphoton signal events from $K^{\pm}\to\pi^{\pm}a$. In shaded gray we show constraints from NA62 [42] for invisible kaon decays, SN1987A [43, 44], and cosmology [45]. In the shaded region labeled “Existing Constraints” we include bounds from invisible kaon decays from E787 and E949 [43], electron beam dump [46, 47, 48], CHARM [49], visible kaon decays [50], $B$ meson decays [39], and LHC di-jet searches [51].
@@ VI Conclusion
In this work, we explored the sensitivity of DUNE ND to heavy axions and millicharged particles produced through charged pion scattering. We found that the expected flux of both ALPs and MCP produced from charged pion scattering was greater than previously-considered production modes meson for some masses in the MeV–GeV range. With this increased flux, we were able to claim a conservative increase in sensitivity of DUNE ND to both ALPs and MCPs. The increase in sensitivity to MCP was modest, but combined with the additional production mechanisms can increase the sensitivity by approximately a factor of 2 at masses above 400 MeV. The increase in sensitivity to ALPs is similar, and combined with the strong projected constraints from kaon decays in Ref. [5], can likely fully probe remaining parameter space for $m_{a}\lesssim 250\ {\rm MeV}$ between the recent NA62 limits and the SN1987A constraint.
@@
Finally, in light of the fact that charged pion scattering dominates the Drell-Yan and gluon fusion production modes for MCPs and axions, respectively, there are likely additional gains in sensitivity to both models from a complete treatment of the high-momentum-transfer regime which interpolates between the chiral Lagrangian and perturbative QCD. Such a treatment would involve, for example, quasi-elastic scattering processes and target mass corrections, and for ALPs would consider the scattering and mixing processes in the same theoretical framework. This study would be highly synergistic with the DUNE neutrino physics program where the intermediate-energy regime is crucial for understanding neutrino-nucleus cross sections [13]. Further improvements in the theoretical framework for GeV-scale neutrino cross sections at DUNE would therefore translate directly into improved new physics sensitivity at DUNE, and we look forward to such collaborations involving lattice QCD, BSM phenomenology, and neutrino physics to maximize the science potential of DUNE.
@@ Acknowledgments
We thank Joshua Berger, Roni Harnik, Kevin Kelly, Zhen Liu, and Yu-Dai Tsai for helpful discussions. We also thank Insung Hwang for their work collecting constraints for MCP detection. This work is supported in part by DOE grant DE-SC0015655.

### v2

@@ IV.2 MCP Detector Signatures and Backgrounds
The proposed MCP detector design is based on that of MilliQan Haas et al. (2015); Ball et al. (2016). The signature for detecting a MCP is three soft ionizations through the scintillator arrays that are collinear with the beam line and target, detectable through the collection of photoelectrons by PMTs. Following Ref. Kelly and Tsai (2019), we also require that all three soft ionizations occur within a small 15 nanosecond window to further reduce background.
There are two types of background that would look like additional soft ionizations throughout the scintillators: detector-related background and beam-related background. To make a direct comparison with Ref. Kelly and Tsai (2019), we make identical assumptions about the backgrounds, yielding an estimate of about 600 events per year; we refer the reader to Ref. Kelly and Tsai (2019) for more details about the background processes.
@@ IV.3 MCP Sensitivity at Dune ND
In the FerMINI design Kelly and Tsai (2019); Patrignani et al. (2016), a minimum-ionizing particle with charge $|Q|=1$ would produce approximately $10^{6}$ photoelectrons (PE) in a 1 m plastic scintillator with a density of $\rho_{\rm scint.}=1$ g/cm3. Thus MCPs with charge $Q=e\epsilon$ would produce PEs that scale as $\epsilon^{2}$. We can then quantify the probability of a MCP producing at least one PE in all three scintillator stacks as
@@
where $\overline{N}_{\rm PE}$ denotes the average number of photoelectrons produced from soft ionization. We use the approximation of $\overline{N}_{\rm PE}(\epsilon)=\left(\frac{\epsilon}{\zeta}\right)^{2}$ where $\zeta=2\times 10^{-3}$. This value of $\zeta$ corresponds to a Saint-Gobain BC-408 plastic scintillator Kelly and Tsai (2019); Ball et al. (2016). Given these choices, such an experiment could be sensitive to $\epsilon$ of order $10^{-4}$.
@@
Assuming that the scintillators can detect values as low as $\epsilon=10^{-4}$, we compute the projected 95% CL limit, i.e. $N_{S}/\sqrt{N_{B}}=2$ where $N_{B}$ is the number of background events, for the DUNE ND setup to MCP from charged pion scattering. The results are shown in Fig. 6 along with constraints and projections from other beam-dump, collider, and direct detection experiments. We also show the projected limit from neutral meson decay and Drell-Yan scattering at DUNE, following Fig. 1 of Ref. Kelly and Tsai (2019) for the total number of MCPs that reach the target, and then applying Eq. (14) in order to model the scintillator response. We find that we reproduce the sensitivity in Ref. Kelly and Tsai (2019) to within about a factor of 2, with any differences likely due to slightly different treatments of the angular acceptance Kelly , which lets us directly assess the increase in sensitivity from charged pion scattering. Overall, the charged pion production channel (which is irreducible in this model) can offer additional sensitivity to detecting MCPs for masses $1.5\ {\rm GeV}\lesssim m_{\chi}\lesssim 3\ {\rm GeV}$.
Figure 6: 95% CL sensitivity to MCP detection at DUNE ND with $10^{21}$ POT (about one year of beam time). The charged pion production channel is shown in purple, and the red-orange curve labeled “Neutral Meson + Drell Yan” is our reproduction of the projected sensitivity from Ref. Kelly and Tsai (2019). The milliQan HL-LHC projected sensitivity Haas et al. (2015); Ball et al. (2016) is plotted in teal, and the LANSCE-mQ at ER1 Tsai et al. (2024) projected sensitivity, using $5.9\times 10^{22}$ POT (about 3 years of beam time), is plotted in blue. Shaded regions are existing limits from SLAC (Prinz (2001), peach), SENSEI (Abramoff et al. (2019), yellow), BEBC (Marocco and Sarkar (2021), light pink), Charm II (De Winter et al. (1989), light green), LSND (Magill et al. (2019), olive green), MiniBooNE (Magill et al. (2019), gray), ArgoNeuT (Acciarri et al. (2020), light blue), and colliders (Davidson et al. (2000), burgundy).
@@ V Heavy Axion-like Particles
The DUNE ND is an excellent experiment to look for long-lived exotic particles at MeV-GeV mass scales. One such particle that we can search for at DUNE ND is a heavy axion coupled to gluons at low energies. This particle may be a generic ALP, but in some models with additional gauge groups, such a heavy axion may in fact be the QCD axion which solves the strong CP problem Hook et al. (2020). In this section, we will be agnostic as to the UV interpretation of such a particle, and simply use this model as a benchmark to make a direct comparison with the model of Ref. Kelly et al. (2021). As we have done with MCPs in Sec. IV, we will demonstrate the additional sensitivity gained from charged pion production, which is an irreducible production channel in this model.
As in Ref. Kelly et al. (2021), we consider a low-energy ALP model where at scales $\mu=\mathcal{O}(1)$ GeV the ALP is only coupled to gluons,
@@
Here, $f_{a}$ is the axion decay constant, $\alpha_{s}$ is the strong fine structure constant, and $G^{a}_{\mu\nu}$ is the gluon field strength tensor. Following Di Luzio et al. (2020); Chang and Choi (1993); Choi et al. (2022), the axion couplings to nucleons and pions in the chiral Lagrangian can be derived by performing a chiral rotation to shift the axion field into the quark mass matrix. This rotation is arbitrary, but in the conventional parameterization which eliminates mass mixing between the axion and the $\pi^{0}$, we obtain
@@
As in Sec. IV, $f_{\pi}=93$ MeV is the pion decay constant, $g_{A}=1.27$ is the nucleon axial coupling, and $g_{0}=0.52$ is another axial coupling extracted from lattice QCD simulations and evaluated at an $\overline{\rm MS}$ scale $\mu=2$ GeV Grilli di Cortona et al. (2016). Note that the chiral rotation which generates these couplings will also generate an axion-photon coupling; in the spirit of this “gluon dominance” model, we will ignore production from this operator and only use it to permit decays $a\to\gamma\gamma$ for $m_{a}<3m_{\pi}$ (see Sec. V.2 below). Furthermore, we note that axion-neutral meson kinetic mixing will also be present in the chiral Lagrangian. In order to compare the parametric dependence of scattering processes to previously-considered mixing processes, we are neglecting diagrams which involve kinetic mixing insertions, and we leave to future work a full analysis which simultaneously treats mixing and scattering in order to capture any potential interference effects.
@@ V.1 ALP Flux
Figure 8: The cross section for the scattering process $\pi N\to\pi Na$ as a function of incoming pion energy $E_{\pi}$ for range of axion masses and $f_{a}=10^{4}$ GeV. The cross section scales with increasing pion energy due to the derivative couplings in the chiral Lagrangian. The cross section also has a weak dependence on $m_{a}$, which is only significant at low pion energies.
@@
We can understand this enhancement through a dimensional analysis estimate. The number of axions produced through neutral pion mixing is Kelly et al. (2021)
@@
This is reasonably close to the $\sim 10^{6}$ enhancement seen in Fig. 9, which accounts for all of the order-1 coupling factors, the true pion energy distribution, and the detector’s geometric acceptance.
@@ V.2 ALP Detector Signatures and Backgrounds
The axions that are created within the angular acceptance of ArgonCube and MPD at DUNE ND can be detected by their decay into SM particles. For the axion masses we are interested in, the axion can decay into photons or hadrons. Even in the gluon-dominated model, axions are able to decay to photons through mixing with neutral mesons (indeed, this was the primary production mechanism of Ref. Kelly et al. (2021)). For $m_{a}<3m_{\pi}$, the axion will primarily decay into two photons, $a\to\gamma\gamma$. The decay width is given by Aloni et al. (2019)
@@
For $3m_{\pi}<m_{a}\lesssim 1$ GeV, the axion decays to primarily to hadrons through $a\to 3\pi$ and $a\to\pi\pi\gamma$ Aloni et al. (2019); Kelly et al. (2021).
We consider the case where the axion decays within the liquid and gaseous argon detectors at DUNE ND, between 574 and 584 m from the target, with an angular acceptance of $\sim$ 3 mrad (see Fig. 1). Approximately 1% of the total number of axions produced are detected due to the small angular acceptance. These axions must also be highly boosted in order to travel 574 m and be within the angular resolution of the detectors. As a result, the decay products are also highly boosted compared to background events Kelly et al. (2021). Much of the background in both detectors comes from neutrino-related scattering processes in the argon tanks Berryman et al. (2020); Kelly et al. (2021). Both the diphoton and hadronic backgrounds would have low energies and would result in a roughly isotropic signature. Since our axion signal is highly boosted, we can easily distinguish signal from background events Kelly et al. (2021).
@@ V.3 ALP Sensitivity at DUNE ND
Combining the expected signal from $a\to\gamma\gamma$ and $a\to$ hadrons, we can determine the axion parameter space to which DUNE ND is sensitive. We calculate the number of signal events according to the methods described in Sec. III.3, using the axion decay length $c\tau_{a}$ from Refs. Kelly et al. (2021); Blinov et al. (2022). The number of axion signal events is then
@@
where $N_{a}=A\Phi_{a}$ with $\Phi_{a}$ in (19), and $\text{BR}(a\to X)$ is the branching ratio of the axion decaying to either photons or hadrons Aloni et al. (2019); Blinov et al. (2022).
In Figure 10, we show the sensitivity of DUNE ND to 3 or more signal events in 10 years of data collection. The increased production rate from charged pion scattering away from meson resonances improves the projected sensitivity of DUNE in both the long-lifetime (small $1/f_{a}$) and short-lifetime (large $1/f_{a}$) regimes. Because we only calculated $N_{a}$ up to the chiral cutoff $t_{n},t_{\pi}<(4\pi f_{\pi})^{2}$, we do not find additional sensitivity beyond the largest $m_{a}$ considered in Ref. Kelly et al. (2021). For axions below the kaon mass, the sensitivity from charged pion scattering alone is within an order-1 factor of axions produced from kaon decay, recently argued in Ref. Berger and Putnam (2024) to be the dominant production channel in this mass range. Including charged pion scattering beyond the first interaction length of the dump could yield additional sensitivity, possibly making these two production modes comparable.
Figure 10: The sensitivity of DUNE ND to 3 or more heavy ALP signal events over 10 years of data collection from charged pion scattering (purple) along with other projections and constraints. We show projections for DUNE ND (solid) along with FASER Ariga et al. (2019) (green dot-dashed). The projection for an axion signal from neutral meson and gluon-gluon fusion production are shown in blue Kelly et al. (2021). The pink curve, adapted from Ref. Berger and Putnam (2024) and rescaled to the same event rate as the other curves, shows diphoton signal events from $K^{\pm}\to\pi^{\pm}a$. In shaded gray we show constraints from NA62 Cortina Gil et al. (2021) for invisible kaon decays, SN1987A Ertas and Kahlhoefer (2020); Chang et al. (2018), and cosmology Depta et al. (2020). In the shaded region labeled “Existing Constraints” we include bounds from invisible kaon decays from E787 and E949 Ertas and Kahlhoefer (2020), electron beam dump Döbrich et al. (2016); Dolan et al. (2017); Banerjee et al. (2020), CHARM Bergsma et al. (1985), visible kaon decays Gori et al. (2020), $B$ meson decays Aloni et al. (2019), and LHC di-jet searches Mariotti et al. (2018).
@@ VI Conclusion
In this work, we explored the sensitivity of DUNE ND to heavy axions and millicharged particles produced through charged pion scattering. We found that the expected flux of both ALPs and MCP produced from charged pion scattering was greater than previously-considered production modes meson for some masses in the MeV–GeV range. With this increased flux, we were able to claim a conservative increase in sensitivity of DUNE ND to both ALPs and MCPs. The increase in sensitivity to MCP was modest, but combined with the additional production mechanisms can modestly increase the sensitivity at masses above 2 GeV. The increase in sensitivity to ALPs is similar, and combined with the strong projected constraints from kaon decays in Ref. Berger and Putnam (2024), can likely fully probe remaining parameter space for $m_{a}\lesssim 250\ {\rm MeV}$ between the recent NA62 limits and the SN1987A constraint.
@@
Finally, it is reassuring that our chiral perturbation theory calculation closely matches the parton-level Drell-Yan or gluon fusion calculations for BSM particle masses around the chiral cutoff, but there are likely additional gains in sensitivity to both MCPs and axions from a complete treatment of the high-momentum-transfer regime which interpolates between the chiral Lagrangian and perturbative QCD. Such a treatment would involve, for example, quasi-elastic scattering processes and target mass corrections, and for ALPs would consider the scattering and mixing processes in the same theoretical framework. This study would be highly synergistic with the DUNE neutrino physics program where the intermediate-energy regime is crucial for understanding neutrino-nucleus cross sections Ruso et al. (2022). Further improvements in the theoretical framework for GeV-scale neutrino cross sections at DUNE would therefore translate directly into improved new physics sensitivity at DUNE, and we look forward to such collaborations involving lattice QCD, BSM phenomenology, and neutrino physics to maximize the science potential of DUNE.
@@ Acknowledgments
We thank Joshua Berger, Roni Harnik, Kevin Kelly, Zhen Liu, and Yu-Dai Tsai for helpful discussions. We also thank Insung Hwang for their work collecting constraints for MCP detection. We thank the CETUP* 2024 conference for facilitating discussions related to this work. This work is supported in part by DOE grant DE-SC0015655.